# ICONS: Influence Consensus for Vision-Language Data Selection

### v3

@@ I.2 Specialists & Generalist
We visualize the most influential top three examples across specialists (figs. 9, 10, 11, 12, 13, 14, 15, 16, 17 and 18) and the generalist selection (Fig. 19), along with samples from their corresponding tasks. Notably, the selected high-influence examples by specialists show strong task-specific characteristics both structurally and contextually - they mirror the key attributes of their target tasks in terms of question structure, reasoning patterns, and required visual-language understanding capabilities. Furthermore, the visualization of top influential examples reveals distinct patterns in what makes training samples valuable for different vision-language tasks. VQAv2, GQA, and SQA-I specialists favor multi-turn Q&A scenarios that test both visual comprehension and contextual understanding, while TextVQA, POPE, and MME specialists emphasize text recognition, object verification, and spatial relationships respectively. MMBench-EN and MMBench-CN show consistent patterns despite language differences, focusing on clear, unambiguous scenes that translate well. The LLaVA-W Bench specialist prioritizes examples requiring detailed explanations and multi-step reasoning, and the answers are generally longer. The generalist model values diverse scenarios that combine multiple skills simultaneously. Common characteristics that make these examples particularly valuable include multi-turn interactions, clear visual elements, factual and inferential reasoning, cross-modal interaction, and the ability to test multiple capabilities within a single example. This suggests that the most effective training samples are those that combine multiple types of reasoning while maintaining clear, unambiguous ground truth that can be consistently learned across tasks.
@@
Figure 9: VQAv2. Top-left: A sample from VQAv2 vqav2 . Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 10: GQA. Top-left: A sample from GQA gqa . Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 11: VizWiz. Top-left: A sample from VizWiz vizwiz . Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 12: SQA. Top-left: A sample from SQA sqa . Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 13: TextVQA. Top-left: A sample from TextVQA textvqa . Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 14: Pope. Top-left: A sample from Pope pope . Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 15: MME. Top-left: A sample from MME mme . Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 16: MMBench (en). Top-left: A sample from MMBench (en) mmbench . Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 17: MMBench (cn). Top-left: A sample from MMBench (cn) mmbench . Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 18: LLaVA-W Bench. Top-left: A sample from LLaVA-W Bench llava . Remaining panels show top three influential samples selected using the specialist influence ranking step.

### v4

@@ F.2 Specialists & Generalist
We visualize the most influential top three examples across specialists (Figs.˜12, 13, 14, 15, 16, 17, 18, 19, 20 and 21) and the generalist selection (Fig. 22), along with samples from their corresponding tasks. Notably, the selected high-influence examples by specialists show strong task-specific characteristics both structurally and contextually - they mirror the key attributes of their target tasks in terms of question structure, reasoning patterns, and required visual-language understanding capabilities. Furthermore, the visualization of top influential examples reveals distinct patterns in what makes training samples valuable for different vision-language tasks. VQAv2, GQA, and SQA-I specialists favor multi-turn Q&A scenarios that test both visual comprehension and contextual understanding, while TextVQA, POPE, and MME specialists emphasize text recognition, object verification, and spatial relationships respectively. MMBench-EN and MMBench-CN show consistent patterns despite language differences, focusing on clear, unambiguous scenes that translate well. The LLaVA-W Bench specialist prioritizes examples requiring detailed explanations and multi-step reasoning, and the answers are generally longer. The generalist model values diverse scenarios that combine multiple skills simultaneously. Common characteristics that make these examples particularly valuable include multi-turn interactions, clear visual elements, factual and inferential reasoning, cross-modal interaction, and the ability to test multiple capabilities within a single example. This suggests that the most effective training samples are those that combine multiple types of reasoning while maintaining clear, unambiguous ground truth that can be consistently learned across tasks.
@@
Figure 12: VQAv2. Top-left: A sample from VQAv2 [vqav2]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 13: GQA. Top-left: A sample from GQA [gqa]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 14: VizWiz. Top-left: A sample from VizWiz [vizwiz]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 15: SQA. Top-left: A sample from SQA [sqa]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 16: TextVQA. Top-left: A sample from TextVQA [textvqa]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 17: Pope. Top-left: A sample from Pope [pope]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 18: MME. Top-left: A sample from MME [mme]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 19: MMBench (en). Top-left: A sample from MMBench (en) [mmbench]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 20: MMBench (cn). Top-left: A sample from MMBench (cn) [mmbench]. Remaining panels show top three influential samples selected using the specialist influence ranking step.
@@
Figure 21: LLaVA-W Bench. Top-left: A sample from LLaVA-W Bench [llava]. Remaining panels show top three influential samples selected using the specialist influence ranking step.